this problem is so difficult can someone please help me out :) thanks
A circular cone having an altitude of 9m is divided into two segments having the same vertex . if the smaller altitude is 6m , find the ratio of the volume of the small cone to the big cone

Answer by stanbon(75887)

A circular cone having an altitude of 9m is divided into two segments having the same vertex . if the smaller altitude is 6m , find the ratio of the volume of the small cone to the big cone
-----
Note: The ratio of the radii of the bases of the two cones is 9:6 or 3/2
---------------------------------------------------
Volume of big cone = (1/3)pi*r^2*9
----
Volume of smaller cone = (1/3)pi(2/3 r)^2*6 = (1/3)pi*r^2[(4/9)*6]
= (1/3)pi*r^2[8/3)
-----
Ans: ratio of volumes = [Vol of small]/[Vol of bigger]
= (8/3)/9 = 8/27 = 8:27
=============
Cheers,
Stan H.